OpenStudy (anonymous):

(m^1/4 n^1/2)^2 (m^2 n^3)^1/2 I need help in solving this problem. Simplifying Radicals by simplify and compare equivalent expressions. Please help.

OpenStudy (anonymous):

(m^1/4 n^1/2)^2 (m^2 n^3)^1/2 (√m2 * 4√n)*(m^2 *n^3)^1/2 Simplify the expression with positive exponents (√m2 * √n)*( √m^2 * √n^3) Product role m* √n *m* √n^3 Multiply each side m*n * √n* √n^3 m^2 * 1√n^2 Multiply exponents m^2 n^2 Positive Exponent Finale answer But my professor said it is incorrect. Can someone tell what step I did wrong please

OpenStudy (anonymous):

Yes that is the problem that I am having trouble figuring out properly.

OpenStudy (anonymous):

I am not sure what step is next.

OpenStudy (cggurumanjunath):

http://prntscr.com/2m26ca

OpenStudy (cggurumanjunath):

http://sketchtoy.com/58520216

OpenStudy (cggurumanjunath):

go through above two links .

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(m^1/4)^2= √m (n^1/2)^2= n (m^2)^1/2= √m^2 (n^3)^1/2=√n^3

OpenStudy (anonymous):

Are those correct so far.

OpenStudy (cggurumanjunath):

yes

OpenStudy (cggurumanjunath):

(m^2)^1/2= √m^2=m

OpenStudy (cggurumanjunath):

(n^3)^1/2=√n^3=n^(3/2)

OpenStudy (cggurumanjunath):

http://sketchtoy.com/58520256===final answer !

OpenStudy (anonymous):

n^(3/2) * n m*√m

OpenStudy (anonymous):

okay I guess I was wrong again then.

OpenStudy (anonymous):

This is totally confusing to me.

OpenStudy (anonymous):

Thank you for all your help. I need to review this a lot more on how you did the steps.

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

tag me immediately when u are stuck at any step. I am always there to help you anytime.

OpenStudy (anonymous):

Thank you that is very nice of you

OpenStudy (cggurumanjunath):

welcome.